# Strategies for implementing quantum error correction in molecular rotation

### v1

@@ A.2 Einstein coefficients and decay rates
$\Gamma^{\textrm{SD}}(J,m,\delta m)=A_{ij}\\ \begin{split}&=\frac{\omega_{ij}^{3}|d_{ij}|^{2}}{3\pi\epsilon_{0}\hbar c^{3}}\\ &=\gamma^{\textrm{SD}}_{J}\bigg{(}\frac{4\pi}{3}\bigg{)}|c^{J}(J-1,m+\delta m,1,\delta m)|^{2}\,,\end{split}$ (49)
@@
$\gamma_{J}^{\textrm{SD}}=\frac{8d^{2}B_{\textrm{R}}J^{3}}{3\pi\epsilon_{0}\hbar^{4}c^{3}}\,,$ (50)
where we assume the frequency differences between the sublevels in each manifold are small such that the rate is only associated with the difference in frequency $\omega_{ij}=\omega_{J,J-1}=2B_{\textrm{R}}J/\hbar$ between the centroids of the rotational manifolds $J$ and $J-1$.
@@
$\begin{split}\hat{\epsilon}_{\delta m}\cdot\hat{d}_{ij}&=d\langle i|\hat{\epsilon}_{\delta m}\cdot\hat{r}|j\rangle\\ &=d\,\sqrt{\frac{4\pi}{3}}\,c^{J_{1}}(J_{2},m_{2},1,\delta m)\\ &=d\,\sqrt{\frac{4\pi}{3}}\iint Y^{*}_{J_{2},m_{2}}(\theta,\phi)Y_{1,\delta m}(\theta,\phi)\\ &\qquad Y_{J_{1},m_{1}}(\theta,\phi)\ \sin{\theta}\ \mathrm{d}\theta\ \mathrm{d}\phi\\ &=d\,\sqrt{(2J_{1}+1)(2J_{2}+1)}\,(-1)^{m_{2}}\\ &\qquad\begin{pmatrix}J_{2}&1&J_{1}\\ -m_{2}&\delta m&m_{1}\end{pmatrix}\begin{pmatrix}J_{2}&1&J_{1}\\ 0&0&0\end{pmatrix},\end{split}$ (52)
@@
$\Gamma^{\textrm{BBR}}(J,\delta m)=B_{ij}\rho(\omega_{ij},T)\\ \begin{split}&=\frac{\omega_{ij}^{3}|d_{ij}|^{2}}{3\pi\epsilon_{0}\hbar c^{3}}\frac{1}{e^{\hbar\omega_{ij}/k_{\textrm{B}}T}-1}\\ &=\gamma^{\textrm{BBR}}_{J}\bigg{(}\frac{4\pi}{3}\bigg{)}|c^{J}(J-1,m+\delta m,1,\delta m)|^{2}\,,\end{split}$ (53)
@@
Interactions with the environment via electric dipole transitions caused by an emisison event due to spontaneous decay (SD) are given by
@@
$\Gamma^{\textrm{SD}}(J,m,\delta m)=A_{ij}\\ =\gamma^{\textrm{SD}}_{J}\bigg{(}\frac{4\pi}{3}\bigg{)}|c^{J}(J-1,m+\delta m,1,\delta m)|^{2}\,,$ (58)
@@
Appendix B Sequential QEC procedures
B.1 Effect of an error on rotational state populations
@@ Appendix C Dissipative QEC
$\frac{\textrm{d}\hat{\rho}(t)}{\textrm{d}t}=-\frac{i}{\hbar}\big{[}\hat{H},\hat{\rho}(t)\big{]}\\ +\sum_{n}\frac{1}{2}\big{[}2\hat{C}_{n}\hat{\rho}(t)\hat{C}_{n}^{\dagger}-\hat{\rho}(t)\hat{C}_{n}^{\dagger}\hat{C}_{n}-\hat{C}_{n}^{\dagger}\hat{C}_{n}\hat{\rho}(t)\big{]}$ (68)
@@ C.1 Dissipative repumping of rotational manifold
To demonstrate the performance of the $J$-repumping scheme, we simulate the evolution of the system without any coupling to BBR. We consider the initial state $\hat{\rho}_{\uparrow}(\hat{\rho}_{\downarrow})=\left|\uparrow\right\rangle\left\langle\uparrow\right|\big{(}\left|\downarrow\right\rangle\left\langle\downarrow\right|\big{)}$, where $\left|\uparrow\right\rangle(\left|\downarrow\right\rangle)=\hat{T}\big{(}7,m,+1(-1),0\big{)}\,|\overline{0}\rangle$ which represents a rotational ladder operator applied to the $|\overline{0}\rangle$ codeword. Upon application of the $J$-repumping scheme with the same parameters as defined in Sec. III.3, the density operator is brought back toward $\hat{\rho}_{0}=|\overline{0}\rangle\langle\overline{0}|$. The expectation value $\langle J\rangle$ returns to $J_{\textrm{C}}=7$ for either initial error state as expected, shown in Fig. 7.
Figure 7: Evolution of the expectation value of $J$ for two different initial error states using the dissipative $J$-repumping scheme.
The evolution of the physical fidelity $F\big{(}\hat{\rho}_{0},\hat{\rho}(t)\big{)}=\textrm{tr}\big{(}\hat{\rho}(t)\hat{\rho}_{0}\big{)}$, after tracing out the motional modes, is shown in Fig. 8. Upon application of the $J$-repumping scheme with the same parameters as defined in Sec. III.3, the density is brought back toward $\hat{\rho}_{0}$. At $t=0.05\,\Gamma_{\textrm{C}}^{-1}$, the physical fidelities reach 0.974 for an initial state $\hat{\rho}_{\downarrow}$ and 0.987 for an initial state $\hat{\rho}_{\uparrow}$. The downward correction of $\hat{\rho}_{\uparrow}$ from $J=8$ reaches a higher fidelity than the upward correction of $\hat{\rho}_{\downarrow}$ from $J=6$ in this time due to the slightly different effective repumping rates arising from the different Slater integrals for each coupling.
Figure 8: Evolution of the physical fidelity $F\big{(}\hat{\rho}_{0},\hat{\rho}(t)\big{)}$ after tracing out the motional modes for two different initial error states using the dissipative $J$-repumping scheme.
@@ C.2 Zeeman dissipative error correction
To demonstrate the performance of the full DEC scheme with rotational manifold repumping and Zeeman DEC combined, we can simulate without any coupling to BBR. We consider the initial states $\hat{\rho}_{\leftarrow}(\hat{\rho}_{\rightarrow})=\left|\leftarrow\rangle\langle\leftarrow\right|\big{(}\left|\rightarrow\rangle\langle\rightarrow\right|\big{)}$, where $\left|\leftarrow\right\rangle(\left|\rightarrow\right\rangle)=\hat{T}\big{(}7,m,0,-1(+1)\big{)}\left|\overline{0}\right\rangle$. This represents a ladder operator applied to the $|\overline{0}\rangle$ codeword followed by perfect repumping back to $J_{\textrm{C}}=7$. The expectation value $\langle J\rangle$ is fixed at $J_{\textrm{C}}=7$ as there is no operator which maps out of the code manifold. Upon application of the Zeeman correction scheme with the same parameters as defined in Sec. III.3, the density operator is brought toward $\hat{\rho}_{0}$.
The evolution of the physical fidelity with respect to $\hat{\rho}_{0}$, after tracing out motional modes, is $F\big{(}\hat{\rho}_{0},\hat{\rho}(t)\big{)}$ and is shown in Fig. 9. After a short time, the physical fidelities for both $m$-shifted states symmetrically reach a value of 0.994 and approach arbitrarily close to 1. There is a small oscillation in the fidelity due to the finite cooling rate, as some population is cycled back to the error state on the red sideband before being returned on the blue sideband and cooled further.
Figure 9: Evolution of the physical fidelity $F\big{(}\hat{\rho}_{0},\hat{\rho}(t)\big{)}$ after tracing out the motional modes for two different initial error states using the Zeeman DEC scheme.
@@ C.3 Full dissipative error correction
The comparison of the action of the full DEC scheme to $J$-repumping only to doing nothing can be observed by examining the final state populations after tracing out the motional modes. This is shown for the same parameters defined in Sec. III.3 in Fig. 10. The concatenation of the Zeeman DEC scheme with the $J$-repumping DEC scheme significantly improves the population preserved in $\mathcal{H}_{\textrm{C}}$.
Figure 10: Final states at $t=2\,\Gamma_{\textrm{C}}^{-1}$ for initial state $\hat{\rho}_{0}$ after doing nothing (top), after $J$-repumping only (middle), and after full DEC (bottom) having traced out the motional modes. The populations of each state are indicated with red squares with areas proportional to the populations. DEC rate parameters are the same as those used in the main text. The code manifold $J_{\textrm{C}}=7$ is outlined in blue.

### v2

@@ A.2 Einstein coefficients and decay rates
$\begin{split}\Gamma^{\textrm{SD}}(J,m,\delta m)&=A_{ij}\\ &=\frac{\omega_{ij}^{3}|d_{ij}|^{2}}{3\pi\epsilon_{0}\hbar c^{3}}\\ &=\gamma^{\textrm{SD}}_{J}\bigg{(}\frac{4\pi}{3}\bigg{)}|c^{J}(J-1,m+\delta m,1,\delta m)|^{2}\,,\end{split}$ (49)
@@
$\gamma_{J}^{\textrm{SD}}=\frac{8d^{2}B_{\textrm{R}}^{3}J^{3}}{3\pi\epsilon_{0}\hbar^{4}c^{3}}\,,$ (50)
where we assume the frequency differences between the sublevels in each manifold are small such that the rate is only associated with the frequency difference ${\omega_{ij}=\omega_{J,J-1}=2B_{\textrm{R}}J/\hbar}$ between the centroids of the rotational manifolds $J$ and ${J-1}$.
@@
$\begin{split}\hat{\epsilon}_{\delta m}\cdot\hat{d}_{ij}&=d\langle i|\hat{\epsilon}_{\delta m}\cdot\hat{r}|j\rangle\\ &=d\,\sqrt{\frac{4\pi}{3}}\,c^{J_{1}}(J_{2},m_{2},1,\delta m)\\ &=d\,\sqrt{\frac{4\pi}{3}}\iint Y^{*}_{J_{2},m_{2}}(\theta,\phi)Y_{1,\delta m}(\theta,\phi)Y_{J_{1},m_{1}}(\theta,\phi)\ \sin{\theta}\ \mathrm{d}\theta\ \mathrm{d}\phi\\ &=d\,\sqrt{(2J_{1}+1)(2J_{2}+1)}\,(-1)^{m_{2}}\begin{pmatrix}J_{2}&1&J_{1}\\ -m_{2}&\delta m&m_{1}\end{pmatrix}\begin{pmatrix}J_{2}&1&J_{1}\\ 0&0&0\end{pmatrix},\end{split}$ (52)
@@
$\begin{split}\Gamma^{\textrm{BBR}}(J,\delta m)&=B_{ij}\rho(\omega_{ij},T)\\ &=\frac{\omega_{ij}^{3}|d_{ij}|^{2}}{3\pi\epsilon_{0}\hbar c^{3}}\frac{1}{e^{\hbar\omega_{ij}/k_{\textrm{B}}T}-1}\\ &=\gamma^{\textrm{BBR}}_{J}\bigg{(}\frac{4\pi}{3}\bigg{)}|c^{J}(J-1,m+\delta m,1,\delta m)|^{2}\,,\end{split}$ (53)
@@
Interactions with the environment via electric dipole transitions caused by an emission event due to spontaneous decay (SD) are given by
@@
$\begin{split}\Gamma^{\textrm{SD}}(J,m,\delta m)&=A_{ij}\\ &=\gamma^{\textrm{SD}}_{J}\bigg{(}\frac{4\pi}{3}\bigg{)}|c^{J}(J-1,m+\delta m,1,\delta m)|^{2}\,,\end{split}$ (58)
@@
Appendix B Effect of an error on rotational state populations
@@ Appendix C Dissipative QEC
$\frac{\textrm{d}\hat{\rho}(t)}{\textrm{d}t}=-\frac{i}{\hbar}\big{[}\hat{H},\hat{\rho}(t)\big{]}+\sum_{n}\frac{1}{2}\big{[}2\hat{C}_{n}\hat{\rho}(t)\hat{C}_{n}^{\dagger}-\hat{\rho}(t)\hat{C}_{n}^{\dagger}\hat{C}_{n}-\hat{C}_{n}^{\dagger}\hat{C}_{n}\hat{\rho}(t)\big{]}$ (68)
@@ C.1 Dissipative repumping of rotational manifold
To demonstrate the performance of the $J$-repumping scheme, we simulate the evolution of the system without any coupling to BBR. We consider the initial state $\hat{\rho}_{\uparrow}(\hat{\rho}_{\downarrow})={\left|\uparrow\right\rangle\left\langle\uparrow\right|\big{(}\left|\downarrow\right\rangle\left\langle\downarrow\right|\big{)}}$, where $\left|\uparrow\right\rangle(\left|\downarrow\right\rangle)={\hat{T}\big{(}7,m,+1(-1),0\big{)}\,|\overline{0}\rangle}$ which represents a rotational ladder operator applied to the ${|\overline{0}\rangle}$ codeword. Upon application of the $J$-repumping scheme with the same parameters as defined in Sec. 3.3, the density operator is brought back toward ${\hat{\rho}_{0}=|\overline{0}\rangle\langle\overline{0}|}$. The expectation value ${\langle J\rangle}$ returns to ${J_{\textrm{C}}=7}$ for either initial error state as expected, shown in Fig. C.3.
The evolution of the physical fidelity ${F\big{(}\hat{\rho}_{0},\hat{\rho}(t)\big{)}=\textrm{tr}\big{(}\hat{\rho}(t)\hat{\rho}_{0}\big{)}}$, after tracing out the motional modes, is shown in Fig. C.3. Upon application of the $J$-repumping scheme with the same parameters as defined in Sec. 3.3, the density is brought back toward $\hat{\rho}_{0}$. At ${t=0.05\,\Gamma_{\textrm{C}}^{-1}}$, the physical fidelities reach 0.974 for an initial state $\hat{\rho}_{\downarrow}$ and 0.987 for an initial state ${\hat{\rho}_{\uparrow}}$. The downward correction of $\hat{\rho}_{\uparrow}$ from ${J=8}$ reaches a higher fidelity than the upward correction of $\hat{\rho}_{\downarrow}$ from ${J=6}$ in this time due to the slightly different effective repumping rates arising from the different Slater integrals for each coupling.
@@ C.2 Zeeman dissipative error correction
To demonstrate the performance of the full DEC scheme with rotational manifold repumping and Zeeman DEC combined, we can simulate without any coupling to BBR. We consider the initial states ${\hat{\rho}_{\leftarrow}(\hat{\rho}_{\rightarrow})=\left|\leftarrow\rangle\langle\leftarrow\right|\big{(}\left|\rightarrow\rangle\langle\rightarrow\right|\big{)}}$, where ${\left|\leftarrow\right\rangle(\left|\rightarrow\right\rangle)=\hat{T}\big{(}7,m,0,-1(+1)\big{)}\left|\overline{0}\right\rangle}$. This represents a ladder operator applied to the ${|\overline{0}\rangle}$ codeword followed by perfect repumping back to ${J_{\textrm{C}}=7}$. The expectation value ${\langle J\rangle}$ is fixed at ${J_{\textrm{C}}=7}$ as there is no operator which maps out of the code manifold. Upon application of the Zeeman correction scheme with the same parameters as defined in Sec. 3.3, the density operator is brought toward $\hat{\rho}_{0}$.
The evolution of the physical fidelity with respect to $\hat{\rho}_{0}$, after tracing out motional modes, is ${F\big{(}\hat{\rho}_{0},\hat{\rho}(t)\big{)}}$ and is shown in Fig. C.3. After a short time, the physical fidelities for both $m$-shifted states symmetrically reach a value of 0.994 and approach arbitrarily close to 1. There is a small oscillation in the fidelity due to the finite cooling rate, as some population is cycled back to the error state on the red sideband before being returned on the blue sideband and cooled further.
@@ C.3 Full dissipative error correction
The comparison of the action of the full DEC scheme to $J$-repumping only to doing nothing can be observed by examining the final state populations after tracing out the motional modes. This is shown for the same parameters defined in Sec. 3.3 in Fig. C.3. The concatenation of the Zeeman DEC scheme with the $J$-repumping DEC scheme significantly improves the population preserved in $\mathcal{H}_{\textrm{C}}$.
[…]
figureEvolution of the expectation value of $J$ for two different initial error states using the dissipative $J$-repumping scheme.
[…]
figureEvolution of the physical fidelity ${F\big{(}\hat{\rho}_{0},\hat{\rho}(t)\big{)}}$ after tracing out the motional modes for two different initial error states using the dissipative $J$-repumping scheme.
[…]
figureEvolution of the physical fidelity ${F\big{(}\hat{\rho}_{0},\hat{\rho}(t)\big{)}}$ after tracing out the motional modes for two different initial error states using the Zeeman DEC scheme.
[…]
figureFinal states at ${t=2\,\Gamma_{\textrm{C}}^{-1}}$ for initial state ${\hat{\rho}_{0}}$ after doing nothing (top), after $J$-repumping only (middle), and after full DEC (bottom) having traced out the motional modes. The populations of each state are indicated with red squares with areas proportional to the populations. DEC rate parameters are the same as those used in the main text. The code manifold ${J_{\textrm{C}}=7}$ is outlined in blue.